OpenStudy (amonoconnor):

"Determine the Infinite Limit: lim[(x+2)/(x+3)] as x approaches 3 from the right." My book says it's negative Infinity, and I don't know how to get this. I got 5/6.... What's the correct process to solve this, to "determine the infinite limit"? Any and all help is greatly appreciated!

OpenStudy (anonymous):

whats the equation u have

OpenStudy (amonoconnor):

lim\[\lim_{x \rightarrow 3+}[(x+2)/(x+3)]\]

OpenStudy (anonymous):

dame idk maybe @cliffordtapp

OpenStudy (thomas5267):

The book probably meant x approaches *-3* from the right since 5/6 is the correct limit.

OpenStudy (amonoconnor):

5/6 is correct? Thank you so much!

OpenStudy (thomas5267):

I believe your book is quite bad since the term infinite limit refers to the limit as x approaches infinity or negative infinity according to my understanding.

OpenStudy (anonymous):

yes he right

OpenStudy (thomas5267):

Since x is approaching a finite number, it would not be a infinite limit although the value of such limit is finite. That is: \[ \lim_{x\to\color{red}{\infty}}f(x)\quad\text{Infinite limit}\\ \lim_{x\to\color{red}{-\infty}}f(x)\quad\text{Infinite limit}\\ \lim_{x\to\color{red}{3}}f(x)=\infty \quad\text{Finite limit}\\ \]

OpenStudy (amonoconnor):

Wait, maybe I really don't know how to do these, but wouldn't it be -1/0 if evaluated with x approaching -3+?

OpenStudy (thomas5267):

EDIT: Not fully awake. Since x is approaching a finite number, it would not be a infinite limit although the value of such limit is *infinite*. That is:

OpenStudy (thomas5267):

One does not simply divide by zero.

OpenStudy (amonoconnor):

Right, obviously... but I'm not "dividing", that's how it would work out to, right?

OpenStudy (amonoconnor):

\[\lim_{x \rightarrow -3+}(x+2)/\lim_{x \rightarrow -3+}(x+3) = (-3+2)/(-3+3) = (-1/0) ?\]

OpenStudy (thomas5267):

Dividing by zero is a meaningless expression in mathematics. If directly evaluating the limit yields \(\dfrac{-1}{0}\) or \(\dfrac{1}{0}\), then you have to find the limit in other ways.

OpenStudy (amonoconnor):

Alright. So... what other ways/avenues?

OpenStudy (thomas5267):

@freckles I will leave this to you. I have no idea how to explain the evaluation of a limit with infinite as its value without using the delta-epsilon thing.

OpenStudy (freckles):

\[\lim_{x \rightarrow -3^+} \frac{x+2}{x+3} \\ \text{ of } -3^+ \text{ means to the right of -3 } \\ \text{ so } x>-3 \\ \text{ now adding 2 on both sides gives } x+2>-3+2=-1 \\ \text{ this means } x+2 \text{ is near } -1 \\ \\ \text{ now adding both sides by 3 we have } x+3>0 \\ \text{ we already know } \lim_{x \rightarrow -3^+}\frac{x+2}{x+3} \text{ is one of the infinities } \\ \text{ and we know the result is } \frac{negative }{positive } \infty \]

OpenStudy (freckles):

if we had \[\lim_{x \rightarrow -3^-}\frac{x+2}{x+3} \\ \text{ the } -3^- \text{ means we are looking to the left of -3 }\] \[\text{ so } x<-3 \\ \text{ which means } x+3<0 \text{ which says } x+3 \text{ is negative } \\ \text{ now adding 2 on both sides } x+2<-1 \\ \text{ so for this one we have } \frac{negative }{negative } \infty \]

OpenStudy (freckles):

and of course neg/pos=neg and neg/neg=pos

OpenStudy (amonoconnor):

I think I understand everything except for where you said that "we already know that (the equation of the limit) is one of the infinities" ...What does that mean? How do you know that?

OpenStudy (freckles):

if you have something/0 where the top something isn't 0 then you know the limit is one of the infinities

OpenStudy (amonoconnor):

That wouldn't mean that the Limit doesn't exist? :/

OpenStudy (freckles):

infinity is not a number so yes you could say the limit does not exist but sometimes the instructor or it is also more meaningful to say what kind of non-existence the limit is if you can

OpenStudy (amonoconnor):

So it's like a limit can be an integer that a function STOPS at, but reaches on a graph/data set nonetheless, but sometimes, the "STOPPING" point is the arbitrarily unreachable value that the graph just gets infinitely close to, in this case -infinity?

OpenStudy (freckles):

the limit being infinity or -negative infinity at x=some finite number is that you have a vertical asymptote there I think something people say the graph blows up there if that makes more sense for example f(x)=1/x looks like:  notice at x=0 we have a vertical asymptote for x>0 the y values are climbing up with non-stop (y values are getting bigger) for x<0 the y values are falling down with non-stop (y values are getting negative big) so we say \[\lim_{x \rightarrow 0^+} \frac{1}{x}= +\infty \text{ or you can just say } \infty \\ \lim_{x \rightarrow 0^-}\frac{1}{x}=-\infty \\ \text{ but yeah it is true in either case you could say } \\ \text{ the limit does not exist } \\ \] you would definitely say \[\lim_{x \rightarrow 0}\frac{1}{x} \text{ does not exist }\]